yea i need to find the slope of the line perpendicular to the graph of each equation. please help thank you
19. y=2x
20. y=-3x
21. y= 7/5x - 2
22.y= -x/5 - 7
23. 2x + 3y = 5
24. y= -8

Answer by Alan3354(69443)

yea i need to find the slope of the line perpendicular to the graph of each equation. please help thank you
19. y=2x
The slope is the coefficient of x, m = 2 for #19
The slope of lines (not THE line, there's an infinite number of them) is the negative inverse.
That's -1/2 for this one.
-------------------
20. y=-3x
21. y= 7/5x - 2
22. y= -x/5 - 7
Do 20 - 22 the same way.
-----------------
23. 2x + 3y = 5
Solve for y, then m is the coefficient of x and do it like the others.
-------------
24. y= -8
In this one, the coefficient of x = 0. The inverse is -1/0 = undefined.